For gcc, the frequency for all loads and stores is 36%. Instruction cache miss rate is 5%. Data cache miss rate is 10%. If a machine has a CPI of 2 without memory stalls and the miss penalty is 40 cycles for all misses, how much faster is a machine with a perfect cache? increase the performance by doubling its clock rate. Since the main memory speed is unlikely to change, assume that the absolute time to handle a cache miss does not change. How much faster will the machine be with the faster clock?
Answer #1

An upvote would be really helpful!

  • Instruction miss rate: 5%
  • Data miss rate: 10%
  • CPI is: 2 (without any memory stalls)
  • Miss penalty: 40 cycles
  • %36 of instructions are load/store

Determine how much faster a machine would run with a perfect cache that never missed.

(I is # of instructions)

Instruction miss cycles = I x 0.05 x 40 = 2 I

Data miss cycles = I x 0.36 x 0.1 x 40 = 1.44 I

Total memory stall cycles = 2 I + 1.44 I = 3.44 I

CPIstall = 2 + 3.44 = 5.44

= 5.44 / 2 = 2.72

Since, how much faster a machine would run with a perfect cache that never missed = 2.72

Now,

Suppose that clock rate of the machine used in the previous example is doubled but the memory speed, cache misses, and miss rate are same. How much faster the machine be with the faster clock?​​​​​​​

Since the clock rate is doubled, new miss penalty will be 2x40=80 clock cycles.

Total memory stall cycles = (0.05 x 80) + 0.36 x (0.10 x 80) = 6.88

CPIfast clock = 2 + 6.88 = 8.88

= (5.44*2)/8.88 = 1.23 times faster